• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name L. Cole

    Log entry time 08:06:24 on April14,2001

    Entry number 60454

    keyword=Owl Shift Summary

    0:00 - Owl shift begins smoothly.

    Run no........................start/stop............................events................................comments
    3529 0:01 - 0:15 1M
    3530 0:16 - 0:31 1M
    3531 0:32 - 0:47 800K ended run; high data rate
    3532 0:49 - 1:05 1M
    3533 1:06 - 1:13 390K ended run; QE measurement
    3534 1:24 - 1:40 1M
    3535 1:41 - 1:57 1M
    3536 1:58 - 2:14 1M
    3537 2:14 - 2:31 1M
    3538 2:32 - 2:48 1M
    3539 2:49 - 3:05 1M
    3540 3:06 - 3:22 1M
    3541 3:22 - 3:38 1M
    3542 3:39 - 3:57 1M
    3543 3:58 - 4:14 1M
    3544 4:14 - 4:43 1M running at 35 uA
    3545 4:44 - 5:21 950K high deadtime (100%)
    3546 N/A 0 junk run
    3547 5:33 - 6:05 1M
    3548 6:05 - 6:38 1M
    3549 6:39 - 7:08 1M
    3550 7:08 - 7:28 636K ended run; CODA problem 3551 N/A 0 junk run
    3552 7:35 - 8:04 1M

    1:14 - MCC phoned and said they had a trip and had to do a QE measurement.

    1:23 - Beam returns.

    4:14 - Changed beam current to 35 uA.

    5:19 - Ended run. Deadtime was 100 %. CODA doesn't want to end run. Restarted CODA.

    7:32 - CODA problem. High deadtime (100%).